# Improved Anonymous Multi Agent Path Finding Algorithm

### 17 Dec 2023 v1

@@ Abstract
We consider an Anonymous Multi-Agent Path-Finding (AMAPF) problem where the set of agents is confined to a graph, a set of goal vertices is given and each of these vertices has to be reached by some agent. The problem is to find an assignment of the goals to the agents as well as the collision-free paths, and we are interested in finding the solution with the optimal makespan. A well-established approach to solve this problem is to reduce it to a special type of a graph search problem, i.e. to the problem of finding a maximum flow on an auxiliary graph induced by the input one. The size of the former graph may be very large and the search on it may become a bottleneck. To this end, we suggest a specific search algorithm that leverages the idea of exploring the search space not through considering separate search states but rather bulks of them simultaneously. That is, we implicitly compress, store and expand bulks of the search states as single states, which results in high reduction in runtime and memory. Empirically, the resultant AMAPF solver demonstrates superior performance compared to the state-of-the-art competitor and is able to solve all publicly available MAPF instances from the well-known MovingAI benchmark in less than 30 seconds.
@@ Introduction
Multi-Agent Path Finding (MAPF) problem is the problem which generally asks to find a set of collision-free paths for a set of agents that operate in a shared environment and have to reach predefined goal locations from the current (start) ones. MAPF has many applications including automated warehouses, autonomous vehicles, video games and is being widely studied in the literature. Depending on the application, many variants of MAPF have been proposed (Stern et al. 2019b) and numerous solutions have been already presented. One variant is the Anonymous MAPF (AMAPF). In AMAPF the agents are interchangeable and each agent may be assigned to any goal, assuming that in the end all goal locations will be reached (in case the number of goal locations is smaller or equal to the number of agents) or each agent will arrive to one goal location (otherwise). This problem naturally arises in such environments where the tasks can be performed by any agent, e.g., think of the identical robots that carry packages/inventory-pods in automated warehouses.
@@
In this work we are interested in solving AMAPF problem optimally w.r.t makespan cost function, which is the arrival time of the last agent. That is, our task is to find a solution where the last agent arrives at its goal location as early as possible. State-of-the-art optimal AMAPF solvers are reduction-based, i.e. the initial problem is reduced to another one and the latter is solved with the off-shelf solvers. In case of AMAPF the reduction is the following. Based on the input graph another one is constructed and then a maximum flow problem on this auxiliary graph is formulated and solved. The latter can be interpreted as finding several paths (subject to certain constraints) on the reduced network. The major bottleneck here is that the size of the network is much larger, both in the number of vertices and edges, compared to the initial AMAPF graph and finding paths on it is burdensome. Moreover, the AMAPF reduction scheme in general assumes that numerous networks may be consecutively constructed (each one being larger than the previous one) and the search should be repeated.
@@
To this end, we present an improved optimal AMAPF solver that follows the reduction-to-the-flow-problem approach but utilizes a novel search method to find paths on the (flow) networks. The crux of this search method is the concept of bulk states and implicit expansions. In brief, instead of generating and expanding numerous search states we compress them into the bulks that form a sequence, exploiting the special structure of the underlying network, and explicitly store in the search tree only the certain representatives of those bulks (while implicitly reasoning about all other states in the bulk). On the theoretical side we show that our search method, which we dub Bulk Search, is complete. On the practical side we compare our improved AMAPF solver, that utilizes Bulk Search, with the state-of-the-art optimal AMAPF solver and show that our algorithm notably scales better to large maps (due to significantly lower number of expansions when finding the paths on the flow networks) and outperforms the competitor on all maps of the well-known MAPF benchmark from (Stern et al. 2019a).
@@
In conventional MAPF formulation (Stern et al. 2019a) a set of agents and their start and goal locations are given as well the specification where each agent starts and what goal it should reach. Even when both the time in discretized into the time steps and workspace is discretized to a graph (which are the two default assumptions in MAPF), obtaining an optimal solution w.r.t. one the most widely-used objectives, e.g. makespan or sum-of-costs, is known to be NP-Hard problem (Yu and LaValle 2013b). Surprisingly, the AMAPF problem, which is a combined problem of both MAPF and goals assignment, can be optimally solved w.r.t. makespan (but not sum-of-costs) in polynomial time (Yu and LaValle 2013a). The seminal method, introduced in (Yu and LaValle 2013a), is based on the reduction of AMAPF to a series of specific graph-search problems, i.e. the problems of finding a maximum flow (Ford Jr and Fulkerson 2015) on a graph of special structure (network) induced by the inpunt MAPF graph. For sum-of-costs objective an adaptation of the seminal MAPF solver, Conflict-Based Search (CBS) (Sharon et al. 2015), was suggested in (Hönig et al. 2018). Indeed, this algorithm is not polynomial. Suboptimal AMAPF was studied in (Okumura and Défago 2022) and several computationally efficient algorithms were proposed in this work which were empirically shown to provide high-quality solutions (however no bound on sub-optimality is theoretically guaranteed). A variant of the AMAPF problem with some additional practically-inspired assumptions, i.e. that the number of goals exceeds the number of agents and thus agents have to move to the new goals upon completing the current ones, was studied in (Nguyen et al. 2017) and solved using the Answer Set Programming (ASP).
@@
More involved variants of AMAPF were studied in (Ma and Koenig 2016; Barták, Ivanová, and Švancara 2021). There it was assumed that the agents are partitioned into the teams (colors) and each team is assigned a set of interchangeable targets (of the same color). In (Ma and Koenig 2016) a combination of CBS and min-cost max-flow algorithm (Ford Jr and Fulkerson 2015) was suggested to solve a problem. In (Barták, Ivanová, and Švancara 2021) several solvers that utilize reduction to SAT were proposed. Indeed, AMAPF can be viewed as a special instantiation of the colored MAPF problem (i.e. when there exists only one team of agents of the same color as all the goals).
@@
Among the other problems that are closely related to AMAPF one can name Lifelong MAPF (LMAPF) (Li et al. 2021) and Multi-agent Pickup and Delivery (MAPD) (Ma et al. 2017). These MAPF variants assume that the agents continuously operate in the environment reaching the specified goals (associated with certain pickup-and-delivery tasks in case of MAPD). However the assignments of goals (tasks) to agents is commonly assumed to be realized by an external procedure and, thus, the assignment sub-problem is not typically considered as a part of a LMAPF/MAPD problem. However the works that consider a combined problem, indeed, exist (Chen et al. 2021; Xu et al. 2022).
@@
Finally, there exist works that study AMAPF in continuous domains, i.e. not assuming that the agents are confined to a given graph but rather allowing them to freely move in the (geometric) workspace (Adler et al. 2015; Solovey and Halperin 2016).
@@
We follow a classical approach (Stern et al. 2019b) to define the problem under investigation – AMAPF. We consider a graph $G=(V,E)$, whose vertices correspond to the locations in the environment and edges – to the transitions between them. $k$ agents are confined to this graph, i.e., initially each agent occupies a (distinct) vertex – $s_{i}$, the start vertex, and at each time step of the discretized timeline it can either wait in its current vertex or move to an adjacent one. The duration of both types of actions (move or wait) is $1$ time step. $k$ goal vertices, $g_{1},...,g_{k}$, are also distinguished and it is assumed that any agent can reach any goal, i.e. there is no pre-defined assignment of agents to the goals.
A plan for an agent, $\pi(s,g)$, is a sequence of (move/wait) actions, s.t. it begins at vertex $s$, ends at vertex $g$ and each action in a plan starts where the the previous action ends. The cost of the plan is the time step by which $g$ is reached. Two plans are said to contain a vertex (similarly – edge) conflict if the agents following them occupy the the same vertex (use the same edge) at the same time step.
The problem now is to find a set of plans $\Pi=\{\pi_{1},...,\pi_{k}\}$, s.t. (1) each pair of plans is conflict-free and (2) all goals are reached. This problem is, in essence, a combination of the assignment problem, where one needs to decide which agent goes to which goal, and the (multi-agent) pathfinding problem, where one need to construct a set of the conflict-free plans.
We consider the following cost objective: $makespan(\Pi)=\max_{i\in{1,...,k}}(cost(\pi_{i}))$, where $cost(\pi)$ is the cost of the individual plan (i.e. the earliest time step when an agent reaches a goal vertex). In this work, we are interested in obtaining makespan-optimal solutions of the problem at hand (AMAPF).
@@
Generally, network flow problem might come in different flavours, see (Ahuja, Magnanti, and Orlin 1995) for an overview. Here we focus on a specific variant of the problem needed for solving AMAPF problems.
A network is a tuple $N=(G,cap,s,g)$, where $G=(V,E)$ is a directed graph, $cap:E\rightarrow Z^{+}$ is the mapping defining the capacities of the edges, $s\in V$ is the source vertex and $g\in V$ is the sink vertex. For a vertex $v\in V$, let $\sigma^{+}(v)$ (resp. $\sigma^{-}(v)$) denote the set of edges of $G$ going to (resp. leaving) $v$. A feasible $s,g$-flow on the network is a mapping $f:E\rightarrow Z^{+}$ that satisfies three types of constraints: edge capacity constraints,
@@
the flow conservation constraints at non terminal vertices,
@@
and the flow conservation constraints at terminal vertices,
$F(f)=\sum_{e\in\sigma^{-}(s)}f(e)=\sum_{e\in\sigma^{+}(g)}f(e)$ (3)
The quantity $F(f)$ is called the value of the flow $f$. Another interpretation of the flow as defined above is that the flow is a set of $s$-$g$ paths in $G$ (possibly overlapping or even duplicating), where each path caries a unit of flow from $s$ to $g$, such that the sum of units passing through any edge does not exceed its capacity.
The standard single-commodity maximum flow problem asks the question: given a network $N$, what is the maximum $F(f)$ that can be pushed through the network? Alternatively, find a set of $s$-$g$ paths that carry the maximum units of flow through the network.
@@
In (Yu and LaValle 2013b), the authors reduced the $T$-steps AMAPF problem, i.e. the one which allows any agent to do at most $T$ actions, to a maximum flow (MF) problem. In specific, it was proved that a $T$-steps AMAPF problem has a solution iff the reduced MF problem has a flow equal to the number of agents. The makespan for an AMAPF instance can therefore be found by finding the smallest $T$ such that $T$-steps AMAPF instance has a solution. We now explain the suggested reduction with a slight modification suggested by (Liu et al. 2019) that simplifies it.
@@
Consider a $T$-steps AMAPF instance with the graph $G=(V,E)$. We first create $12T+$ 1 copies of $V$ and mark them as follows: $0,1,1^{\prime},2,2^{\prime},...,T^{\prime}$ – see Fig. 1. Hereinafter, we will use the term vertices to denote the elements of the original AMAPF graph and the term nodes to denote the elements of the constructed network. We will also call copies $t^{\prime}$ (with apostrophe) and copies $t$ for $t=0,1,..T$ as outer and inner copies, respectively. The copied vertices of the original graph form the nodes of the network. Each node is identified by $(v,h)$, where $h$ stands for the copy, alternatively referred to as height. Higher node means greater copy, and $h^{\prime}$ is higher than $h$. Indeed, the node $(v,h)$ (as well as $(v,h^{\prime})$) corresponds to the state of an agent located in vertex $v$ at time step $h$.
@@
Then, for each edge $e(u,v)$ in the original graph, we connect the nodes $(u,h^{\prime})$ and $(v,h+1)$ for $h>0$, and also connect $(u,0)$ and $(v,1)$. These edges correspond to the move actions and we call them the move edges. Then we add the wait edges that connect the nodes $(u,h^{\prime})$ and $(u,h+1)$ for $h>0$ and the nodes $(u,0)$ and $(u,1)$. Additionally, we add the edges between the nodes $(u,h)$ and the $(u,h^{\prime})$. These edges do not denote to any action but are added to forbid any two $s$-$g$ paths in the network to share any node and therefore to avoid node-collisions. We call them the restriction edges. Finally, we add the source $s$ and sink $g$ nodes and connect $s$ to all nodes $(u,0)$ where $u$ is a start vertex, and connect $g$ to nodes $(v,T^{\prime})$ where $v$ is a goal vertex.
@@
Hereinafter whenever a path in a network is mentioned we mean $s$-$g$ path. The matching between the AMAPF and MF is now straightforward. Each plan for an agent can be matched to a path in the network by matching the agent actions with the move or wait edges, and using the restriction edges to connect between them. In the same way, a path in the network is matched to a plan for an agent where the move/wait edges are matched to move/wait actions. See Fig. 1 for an self-contained example. It was proved that there are no shared nodes in any two paths in the network (that form the solution to the MF problem) which infers that all matched plans have no node-collisions. Edge-collision may happen if two paths pass the edges $((u,h^{\prime})\rightarrow(v,h+1))$ and $((v,h^{\prime})\rightarrow(u,h+1))$ as these two different edges in the network refer to the same edge in the original graph. However, using the approach suggested in (Liu et al. 2019) these collision can be eliminated in the following fashion. Instead of two conflicting agents move to their next vertices they swap plans and continue moving by each others plan. As a result, the agents’ plans can be obtained by finding the maximum number of paths (maximum flow) in the described network.
@@
Figure 1: Example of the flow network (left) for a $T$-steps AMAPF instance (right). Each line in the flow network presents the copies of a single vertex in the original AMAPF graph. The diagonal, solid horizontal and dashed horizontal edges in the network denote move, wait and restriction edges. In this example, the AMAPF instance has two start vertices $A,B$ and two goal vertices $C,D$, so the source node $s$ in the network is connected to the nodes $(A,0),(B,0)$ and the sink $g$ is connected to $(C,T^{\prime}),(D,T^{\prime})$. The example also shows the matching between the plans for the agents and the $s$-$g$ paths in the network (green and yellow edges).
@@
Ford-Fulkerson algorithm (Ford and Fulkerson 1956) was suggested in (Yu and LaValle 2013b) to solve the maximum flow problem. The algorithm is simple, easy to implement and its complexity on the reduced networks is $O(kEV)$ (formulated in (Yu and LaValle 2013b)) , where $E$, $V$ stand for the number of nodes and edges in the original graph. Since in our MF problem, all edges have a capacity of one and the value of the flow is bounded by the number of agents (as each path is to be matched to an agent’s plan), Ford-Fulkerson can be considered as one of the most efficient solvers for our case (see (Cruz-Mejía and Letchford 2023) for more wider details about Maximum Flow algorithms).
We refer the reader to original paper for the detailed description of Ford-Fulkerson and here just briefly describe how the algorithm works specifically for the reduced network. In particular, the following two steps are sequentially repeated. First, we find a path $p$ from $s$ to $g$ in the network. Second, we reverse all edges of $p$. We keep repeating these two steps until no path can be found. The found paths form the maximum flow in the network (in our case, they form the plans of the agents in the original AMAPF problem).
@@
The maximum value of $T$ needed to solve AMAPF problem can be up to $k+V-2$ (as shown by (Yu and LaValle 2013b)). This implies that the network size may be quadratic in the number of vertices of the original graph. Therefore, finding an $s$-$g$ path may become a bottleneck when solving AMAPF instances involving large input graphs. To this end, we propose an algorithm to enhance the search process taking advantage from the specific structure of the reduced network and as a result solving AMAPF problem much faster.
First, we present some definitions to use in the algorithm description. In our algorithm, the search state corresponds to the network node. We will use $n(v,h)$ to denote the search states. Recall that the network node is defined by the vertex in original graph, $v$, and the height (copy) $h$ (higher node means higher copy, and $h^{\prime}$ is higher than $h$). Let us define a connected-sequence as a sequence of nodes with the same vertex in which we can achieve the last node from the first node using only wait and restriction not-reversed edges, and it cannot be extended in neither side (i.e. it has maximum length). An example is depicted in Fig. 2. Initially, for each vertex $v$ of the original graph we have only one connected-sequence $(v,[0,T^{\prime}])$ i.e. the one that starts with the node with height 00 and ends with the node of height $T^{\prime}$ (shown by yellow crossbars). After a path (green path) is found, its edges are reversed (red edges in lower graph). As a result, some connected-sequences disconnect which results in several connected-sequences at the same vertex (see the lower network).
@@
The suggested algorithm is a graph traversal algorithm where the order of the search states in the search set (OPEN) is determined by their heights, i.e. the states with the lower heights are expanded first. The crucial idea of the algorithm is to expand states in bulks while searching. That is, instead of expanding one search state, i.e. generating its successors and marking it as visited (adding it to CLOSED), in each search iteration we (implicitly) expand a bulk of states at once. Such a bulk expansion can be effectively implemented using the introduced notion of the connected sequence, resulting in the reduction of the expansions number, time and memory compared to expanding states individually. Next, we describe how we can form the bulks of states and how the successors can be found compactly and fast.

### 19 Dec 2023 v2

# Improved Anonymous Multi-Agent Path Finding Algorithm

@@ Abstract
We consider the Anonymous Multi-Agent Path-Finding (AMAPF) problem where the agents are confined to a graph, a set of goal vertices is given, and each of these vertices has to be reached by some agent. The problem is to find an assignment of the goals to the agents as well as the collision-free paths, and we seek to find the solution with the minimal makespan. A well-established approach to solving this problem is by reducing it to a special type of graph search problem, i.e. to the problem of finding a maximum flow on an auxiliary graph induced by the input one. The size of the former graph may be very large, and the search on it may become a bottleneck. To this end, we suggest a specific search algorithm that leverages the idea of exploring the search space not through considering separate search states but rather bulks of them simultaneously. That is, we implicitly compress, store, and expand bulks of the search states as single states, reducing the runtime and memory consumption. Empirically, the resultant AMAPF solver demonstrates superior performance compared to the state-of-the-art competitor and is able to solve all publicly available MAPF instances from the well-known MovingAI benchmark in less than 30 seconds.
@@ Introduction
The Multi-Agent Path Finding (MAPF) problem is a problem which generally asks to find a set of collision-free paths for a set of agents that operate in a shared environment and have to reach predefined goal locations from the current (start) ones. MAPF has many applications, including automated warehouses, autonomous vehicles, and video games and is being widely studied in the literature. Depending on the application, many variants of MAPF have been proposed (Stern et al. 2019b) and numerous solutions have been already presented. One variant is the Anonymous MAPF (AMAPF). In AMAPF, the agents are interchangeable and each agent may be assigned to any goal, assuming that, in the end, all goal locations will be reached (in case the number of goal locations is smaller than or equal to the number of agents) or each agent will arrive to one goal location (otherwise). This problem naturally arises in such environments where the tasks can be performed by any agent, e.g. identical robots carrying packages/inventory pods in automated warehouses.
@@
In this work we aim to solve the AMAPF problem optimally w.r.t makespan cost function, which is the arrival time of the last agent. In other words, our task is to find a solution where the last agent arrives at its goal location as early as possible. State-of-the-art optimal AMAPF solvers are reduction-based, i.e. the initial problem is reduced to another one and the latter is solved with an off-shelf solver. In the case of AMAPF, the common reduction is the following. Based on the input graph, another one is constructed. Then a maximum flow problem on this auxiliary graph, called a network, is formulated and solved. The latter can be interpreted as finding several paths (subject to certain constraints) on the reduced network. The major bottleneck here is that the size of the network is much larger, both in the number of vertices and edges, compared to the initial AMAPF graph, and, therefore, finding paths on it is burdensome. Moreover, the AMAPF reduction scheme, in general, assumes that numerous networks may be consecutively constructed (each one being larger than the previous one) and the search should be repeated.
@@
To this end, we present an improved optimal AMAPF solver that follows the reduction-to-the-flow-problem approach while utilizing a novel search method to find paths on the (flow) networks. The crux of our search method is the concept of bulk states and implicit expansions. In brief, instead of generating and expanding numerous search states, we compress them into the bulks that form a sequence, exploiting the special structure of the underlying network, and explicitly store in the search tree only the certain representatives of those bulks (while implicitly reasoning about all other states in the bulk). On the theoretical side, we show that our search method, dubbed Bulk Search, is complete. On the practical side, we compare our improved AMAPF solver that utilizes Bulk Search with the state-of-the-art optimal AMAPF solver and show that our algorithm notably scales better to large maps (due to significantly lower number of expansions when finding the paths on the flow networks) and outperforms the competitor on all maps of the well-known MAPF benchmark from (Stern et al. 2019a).
@@
In a conventional MAPF formulation (Stern et al. 2019a) a set of agents is given as well as the specification where each agent starts and the goal it should reach. Even when both the time is discretized into time steps and workspace is discretized into a graph (which are the two default assumptions in MAPF), obtaining an optimal solution w.r.t. one of the most widely-used objectives, e.g. the makespan or the sum-of-costs, is known to be NP-hard (Yu and LaValle 2013b). Surprisingly, the AMAPF problem, which is a combined problem of both MAPF and goal assignment, can be optimally solved w.r.t. makespan (but not the sum-of-costs) in polynomial time (Yu and LaValle 2013a). The seminal method, introduced in (Yu and LaValle 2013a), is based on the reduction of AMAPF to a series of specific graph-search problems, i.e. the problems of finding a maximum flow (Ford Jr and Fulkerson 2015) on a graph of special structure (network) induced by the input MAPF graph. For the sum-of-costs objective an adaptation of the seminal MAPF solver, Conflict-Based Search (CBS) (Sharon et al. 2015) was suggested in (Hönig et al. 2018). Indeed, this algorithm is not polynomial. Suboptimal AMAPF was studied in (Okumura and Défago 2022); several computationally efficient algorithms were proposed in this work which were empirically shown to provide high-quality solutions. However, no bound on sub-optimality was theoretically guaranteed. A variant of the AMAPF problem with some additional practically inspired assumptions, i.e., that the number of goals exceeds the number of agents and thus agents have to move to the new goals upon completing the current ones, was explored in (Nguyen et al. 2017) and solved using the Answer Set Programming (ASP).
@@
More involved variants of AMAPF were studied in (Ma and Koenig 2016; Barták, Ivanová, and Švancara 2021). It was assumed that the agents are partitioned into the teams (colors) and each team is assigned a set of interchangeable targets (of the same color). In (Ma and Koenig 2016), a combination of CBS and min-cost max-flow algorithm (Ford Jr and Fulkerson 2015) was suggested to solve this Colored MAPF problem. Barták, Ivanová, and Švancara (2021) proposed several solvers that utilize reduction to SAT. Indeed, AMAPF can be viewed as a special instantiation of the Colored MAPF problem (i.e. the one when there exists only a single team of agents of the same color as all the goals).
@@
Among the other problems that are closely related to AMAPF, one can name Lifelong MAPF (LMAPF) (Li et al. 2021) and Multi-agent Pickup and Delivery (MAPD) (Ma et al. 2017). These MAPF variants assume that the agents continuously operate in the environment reaching the specified goals (associated with certain pickup-and-delivery tasks in the case of MAPD). However, the assignments of goals (tasks) to agents is commonly assumed to be realized by an external procedure and, thus, the assignment sub-problem is not typically considered as part of the LMAPF/MAPD problem. Still, there are papers that consider a combined problem (Chen et al. 2021; Xu et al. 2022).
@@
Finally, a body of works studies AMAPF in continuous domains, i.e. not assuming that the agents are confined to a given graph but are rather allowed to freely move in the (geometric) workspace (Adler et al. 2015; Solovey and Halperin 2016).
@@
We follow a classical approach (Stern et al. 2019b) to define the problem under investigation – AMAPF. We consider a graph $G=(V,E)$, whose vertices correspond to the locations in the environment and edges – to the transitions between them. $k$ agents are confined to this graph, i.e., initially each agent occupies a (distinct) vertex – $s_{i}$, the start vertex, and at each time step of the discretized timeline, it can either wait in its current vertex or move to an adjacent one. The duration of both types of actions (move or wait) is $1$ time step. $k$ goal vertices, $g_{1},...,g_{k}$, are also distinguished. It is assumed that any agent can reach any goal, i.e., there is no pre-defined assignment of agents to the goals.
A plan for an agent, $\pi(s,g)$, is a sequence of (move/wait) actions, s.t. it begins at vertex $s$ and ends at vertex $g$; each action in the plan starts where the the previous one ends. The cost of the plan is the time step by which $g$ is reached. Two plans are said to contain a vertex (similarly, an edge) conflict if the agents following them occupy the same vertex (use the same edge) at the same time step.
The problem now is to find a set of plans $\Pi=\{\pi_{1},...,\pi_{k}\}$, s.t. (1) each pair of plans is conflict-free and (2) all goals are reached. Essentially, this problem is a combination of the assignment problem, where one needs to decide which agent goes to which goal, and the (multi-agent) pathfinding problem, where one needs to construct a set of conflict-free plans.
We consider the following cost objective: $makespan(\Pi)=\max_{i\in{1,...,k}}(cost(\pi_{i}))$, where $cost(\pi)$ is the cost of the individual plan (i.e. the earliest time step when the agent reaches a goal vertex and never moves away). In this work, we are interested in obtaining makespan-optimal solutions of the problem at hand (AMAPF).
@@
Generally, network flow problem might come in different flavors; see (Ahuja, Magnanti, and Orlin 1995) for an overview. Here we focus on a specific variant of the problem needed for solving AMAPF problems.
A network is a tuple $N=(G,cap,s,g)$, where $G=(V,E)$ is a directed graph, $cap:E\rightarrow Z^{+}$ is the mapping defining the capacities of the edges, $s\in V$ is the source vertex, and $g\in V$ is the sink vertex. For vertex $v\in V$, let $\sigma^{+}(v)$ (resp. $\sigma^{-}(v)$) denote the set of edges of $G$ going to (resp. leaving) $v$. A feasible $s,g$-flow on the network is mapping $f:E\rightarrow Z^{+}$ that satisfies three types of constraints: edge capacity constraints:
@@
; the flow conservation constraints at non-terminal vertices:
@@
; and the flow conservation constraints at terminal vertices:
$F(f)=\sum_{e\in\sigma^{-}(s)}f(e)=\sum_{e\in\sigma^{+}(g)}f(e).$ (3)
The quantity $F(f)$ is called the value of the flow $f$. Another interpretation of the flow is that the flow is a set of $s$-$g$ paths (possibly overlapping or even duplicating), where each path caries a unit of flow from $s$ to $g$, such that the sum of units passing through any edge does not exceed its capacity.
The standard single-commodity maximum flow problem asks the following question: Given a network $N$, what is the maximum $F(f)$ that can be pushed through the network? Alternatively, find a set of $s$-$g$ paths that carry the maximum units of flow through the network.
@@
In (Yu and LaValle 2013b), the authors reduced the $T$-steps AMAPF problem, i.e. the one that allows any agent to do at most $T$ actions, to a maximum flow (MF) problem. Specifically, it was proven that a $T$-steps AMAPF problem has a solution iff the reduced MF problem has a flow equal to the number of agents. The makespan for an AMAPF instance can therefore be found by finding the smallest $T$ such that $T$-steps AMAPF instance has a solution. We now explain the suggested reduction with a slight modification suggested by (Liu et al. 2019) that simplifies it.
@@
Consider a $T$-steps AMAPF instance with the graph $G=(V,E)$. We first create $12T+$ 1 copies of $V$ and mark them as follows: $0,1,1^{\prime},2,2^{\prime},...,T^{\prime}$; see Fig. 1. Hereinafter, we will use the term vertices to denote the elements of the original AMAPF graph and the term nodes to denote the elements of the constructed network. We will also call copies $t^{\prime}$ (with apostrophe) and copies $t$ for $t=0,1,..T$ as outer and inner copies, respectively. The copied vertices of the original graph form the nodes of the network. Each node is identified by $(v,h)$, where $h$ stands for the copy, alternatively referred to as height. Indeed, the node $(v,h)$ (as well as $(v,h^{\prime})$) corresponds to the state of an agent located in vertex $v$ at time step $h$.
@@
Then, for each edge $e(u,v)$ in the original graph, we connect the nodes $(u,h^{\prime})$ and $(v,h+1)$ for $h>0$, and also connect $(u,0)$ and $(v,1)$. These edges correspond to the move actions, and we call them the move edges. Then, we add the wait edges that connect the nodes $(u,h^{\prime})$ and $(u,h+1)$ for $h>0$ and the nodes $(u,0)$ and $(u,1)$. Additionally, we add the edges between the nodes $(u,h)$ and the $(u,h^{\prime})$. These edges do not denote any action but are added to forbid any two $s$-$g$ paths in the network from sharing any node and, therefore, to avoid node-collisions. We call them the restriction edges. Finally, we add the source $s$ and sink $g$ nodes and connect $s$ to all nodes $(u,0)$ where $u$ is a start vertex, and connect $g$ to nodes $(v,T^{\prime})$ where $v$ is a goal vertex. The capacity of any edge is fixed to be 1.
@@
Hereinafter, whenever a path in a network is mentioned we will mean the $s$-$g$ path. The matching between the AMAPF and MF is now straightforward. Each plan for an agent can be matched to a path in the network by matching the agent actions to the move or wait edges and using the restriction edges to connect between them. Similarly, a path in the network is matched to a plan for an agent where the move/wait edges are matched to move/wait actions. See Fig. 1 for a self-contained example. It was proven that there are no shared nodes in any two paths in the network (that form the solution to the MF problem), which infers that all matched plans have no node-collisions. An edge-collision may happen if two paths pass the edges $((u,h^{\prime})\rightarrow(v,h+1))$ and $((v,h^{\prime})\rightarrow(u,h+1))$ as these two different edges in the network refer to the same edge in the original graph. However, using the approach suggested in (Liu et al. 2019), these collision can be eliminated in the following fashion. Instead of two conflicting agents moving to their next vertices, they swap plans and continue moving by the other’s plan. As a result, the AMAPF solution can be obtained by finding the maximum number of paths (maximum flow) in the described network.
@@
Figure 1: Example of the flow network (left) for a $T$-steps AMAPF instance (right). Each line in the flow network represents the copies of a single vertex in the original AMAPF graph. The diagonal, solid horizontal, and dashed horizontal edges in the network denote move, wait, and restriction edges, respectively. In this example, the AMAPF instance has two start vertices $A,B$ and two goal vertices $C,D$, so the source node $s$ in the network is connected to the nodes $(A,0),(B,0)$ and the sink $g$ is connected to $(C,T^{\prime}),(D,T^{\prime})$. The example also shows the matching between the plans for the agents and the $s$-$g$ paths in the network (green and yellow edges).
@@
Ford-Fulkerson algorithm (Ford and Fulkerson 1956) was suggested in (Yu and LaValle 2013b) to solve the maximum flow problem. The algorithm is simple and easy to implement; its complexity on the reduced networks is $O(kEV)$ (formulated in (Yu and LaValle 2013b)), where $E$, $V$ stand for the number of nodes and edges in the original graph. Since, in our MF problem, all the edges have a capacity of one and the value of the flow is bounded by the number of agents (as each path is to be matched to an agent’s plan), Ford-Fulkerson can be considered as one of the most efficient solvers for our case (see (Cruz-Mejía and Letchford 2023) for more details about Maximum Flow algorithms).
We refer the reader to the original paper for the detailed description of Ford-Fulkerson while briefly describing how the algorithm works specifically for the reduced network. In particular, the following two steps are sequentially repeated. First, we find a path $p$ from $s$ to $g$ in the network. Second, we reverse all edges of $p$. We keep repeating these two steps until no path can be found. The found paths form the maximum flow in the network (in our case, they form the plans of the agents in the original AMAPF problem)  [^1] Arxiv version of this paper contains an illustrative example..
@@
The maximum value of $T$ needed to solve an AMAPF problem can be up to $k+V-2$, as shown by (Yu and LaValle 2013b). This implies that the network size may be quadratic in the number of vertices of the original graph. Therefore, finding an $s$-$g$ path may become a bottleneck when solving AMAPF instances involving large input graphs. To this end, we propose an algorithm that takes advantage of the specific structure of the reduced network and, as a result, is able to solve AMAPF problems much faster.
First, we present some definitions to use in the algorithm description. In our algorithm, the search state corresponds to the network node. We will use $n(v,h)$ to denote the search state. Recall that the network node is defined by the vertex in original graph, $v$, and the height (copy) $h$ (a higher node means a higher copy, and $h^{\prime}$ is higher than $h$). Let us define a connected-sequence as a sequence of nodes with the same vertex in which we can achieve the last node from the first node using only wait and restriction not-reversed edges, and it cannot be extended in either side (i.e., it has the maximum length). An example is shown in Fig. 2. Initially, for each vertex $v$ of the original graph, we have only one connected-sequence $(v,[0,T^{\prime}])$ i.e. the one that starts with the node with height 00 and ends with height $T^{\prime}$ (shown by yellow crossbars). After a path (green path) is found, its edges are reversed (red edges in the lower graph). As a result, some connected-sequences disconnect which leads to several connected-sequences at the same vertex (see the lower network).
@@
The suggested algorithm is a graph traversal algorithm where the order of the search states in OPEN (the set of nodes that are the candidates to be expanded at the next iteration of the algorithm) is determined by their heights, i.e., the states with the lower heights are expanded first. The crucial idea of the algorithm is to expand states in bulks while searching. That is, instead of expanding one search state, i.e. generating its successors and marking it as visited (adding it to CLOSED), in each search iteration, we (implicitly) expand a bulk of states at once. Such bulk expansion can be effectively implemented using the introduced notion of the connected sequence, resulting in the reduction of the expansions number, time and memory compared to expanding states individually. Next, we describe how we can form the bulks of states and how the successors can be found compactly and fast.